A 0.4230 g sample of impure sodium nitrate was heated, converting all the sodium nitrate to .1235 g of sodium nitrate and oxygen gas. Determine the percent of sodium nitrate in the original sample.
1
Expert's answer
2019-03-21T08:10:31-0400

Answer:

2HNO3 = 2HNO2 + O2

m= 85 x 0.1235 / 69= 0.1521

W%= 0.1521 / 0.4230 = 36%